# Estimating the Conformal Prediction Threshold from Noisy Labels

### v1

@@ 4 Experiments
Table 4: Finite sample correction terms $\Delta$ of NACP and ACNL [17] for several datasets and two noise levels, $n$ is the size of the validation set.
Dataset $n$classes NACP ACNL
$\epsilon=0.1$ $\epsilon=0.2$ $\epsilon=0.1$ $\epsilon=0.2$
CIFAR-10 5000 10 0.0351 0.0431 0.0312 0.0594
CIFAR-100 10000 100 0.0248 0.0305 0.0769 0.1634
TinyImagenet 5000 200 0.0351 0.0431 0.1745 0.3819
ImageNet 25000 1000 0.0157 0.0193 0.1936 0.4658
Figure 1: Histograms showing the distribution of (1) $q$ and (2) mean size values across 1000 splits on CIFAR-10 for the rand-APS score and the CP (Oracle), Noisy-CP, and NACP methods.
Figure 2: Mean size and coverage as a function of the noise level using rand-APS for PathMNIST and $0.91\!-\!\alpha=$ 0.9. The calibration set size is $n=8592$ (Table 2).
Table 5: Rand-APS calibration results for $1\!-\!\alpha$ = 0.9 on CIFAR-100 dataset and two noise models. We report the mean and the std over 1000 different splits.
Neighborhood noise Random noise
CP Method size $\downarrow$ coverage (%) size $\downarrow$ coverage (%)
CP (oracle) 6.48 $\pm$ 0.19 90.01 $\pm$ 0.41 6.48 $\pm$ 0.19 90.01 $\pm$ 0.41
Noisy-CP 48.89 $\pm$ 1.13 99.80 $\pm$ 0.04 50.25 $\pm$ 1.37 99.82 $\pm$ 0.04
NR-CP 12.82 $\pm$ 0.36 95.62 $\pm$ 0.21 37.01 $\pm$ 0.88 99.53 $\pm$ 0.06
NACP (w/o $\Delta$) 6.52 $\pm$ 0.22 90.03 $\pm$ 0.47 6.45 $\pm$ 0.30 89.97 $\pm$ 0.57
[…]
Different noise levels and End-to-end CP with noise level estimation. Next, we address one of the basic limitations of our work, that the noise level needs to be known in advance. We show that our approach works well even when we use an estimation fo the noise level. To do so, we combined our NACP with a noise-robust network training that estimated the noise level $\epsilon$ as part of the training [9]. Fig. 2 reports the mean size and coverage results across different noise levels $\epsilon$ and different conformal prediction methods on the PathMNIST dataset. We denote the NACP variant in which $\epsilon$ was estimated by NACP. The results show that as the noise rate increased, the noisy labels corrupted the calibration of the network more, as reflected in a larger mean size and higher coverage, whereas our method, even in the case where $\epsilon$ was estimated achieved much better calibration results which were on par with the result of the noise-free CP.
General noise transition matrix. Finally, we evaluate NACP on two common general noise matrices: Neighborhood noise and Random noise (see details in the supplementary). While existing final sample terms bounds are not effective, in practice NACP (without a finite sample correction) achieves the required coverage guarantee and the average prediction size is similar to the one obtained by the noise-free CP. We observe the same pattern when using uniform noise. This indicates that the current coverage guarantee bounds are too conservative. Table LABEL:generalNoiseResults shows the results on the CIFAR-100 dataset and the rand-APS technique when using NACP without finite sample correction term $\Delta$. Results show a clear dominance of NACP over Noisy-CP and NRCP on the two different noise models, presenting the robustness of NACP across various noise models. ACNL (without the finite sample term) achieves here similar results.
5 Conclusions
We presented a procedure that applies the Conformal Prediction algorithm on a validation set with noisy labels. We first presented our method in the simpler case of a uniform noise model and then extended it to a general noise matrix. We showed that if the noise level is given, we can find the noise-free calibration threshold without access to clean data by using the noisy-label data. We provided a finite-sample coverage guarantee for the case of uniform noise. We showed that in case of uniform noise our method outperforms current noisy CP methods by a large margin, in terms of the average size of the prediction set, while maintaining the required coverage. In all the experiments we conducted the finite sample term was effective and yielded a small prediction set. We showed, however, that even without adding the finite term we obtained the required coverage. This indicates that the current coverage guarantee analysis is too conservative and there is room for future research to improve it. In this study, we focused on noise models that assume the noisy label and the input image are independent, given the true label. In a more general noise model, the label corruption process also depends on the input features.
Appendix A Appendix / supplemental material
A.1 General noise matrices
We define the two common general noise matrices. The Neighborhood noise as:
$P_{i,j}=p(\tilde{y}=j|y=i)=\begin{cases}\xi\hskip 48.36958pt\text{if }i=j\\ 1-\xi\hskip 31.2982pt\text{if }|i-j|=1\text{ and }i\in(1,k)\\ (1-\xi)/2\hskip 14.22636pt\text{if }|i-j|=1\text{ and }i\not\in(1,k)\\ 0\hskip 46.94687pt\text{otherwise}\end{cases}$ (19)
The Random noise is defined as: first, on the diagonal, we have $\xi$. next, for each line (aka $\forall i$) the rest of the values (i.e. $k-1$ items) are sampled from a random distribution $u_{i}$ vector of size $k-1$ and then normalized to sum up to $1-\xi$ to keep the matrix a probability matrix.
$P_{i,j}=p(\tilde{y}=j|y=i)=\begin{cases}\xi\hskip 110.96556pt\text{if }i=j\\ (1-\xi)\cdot\frac{u_{i}[j]}{\sum_{z\neq i}u_{i}[z]}\hskip 36.98866pt\text{otherwise}\end{cases}$ (20)
In our experiments, we set $\xi$ such that the diagonal would be the same as the uniform noise experiments , i.e. $\xi=1-\epsilon+\epsilon/k$ where $\epsilon=0,2$.
A.2 A finite sample term for the case of a general noise matrix
Theorem A.1.
Let $P$ be a general noise matrix. Given $\delta>0$, define $\Delta=\|P^{-1}\|_{\infty}k\sqrt{\frac{\log(2k^{2}/\delta)}{2n}}$ where , $k$ is the number of classes and $n$ is the size of the noisy validation set. Then
$p(\sup\limits_{q}|\hat{F}^{c}(q)-{F}^{c}(q)|>\Delta)<\delta.$
Proof.
From Eq. (18) we have $|\hat{F}^{c}(q)-F^{c}(q)|=|\operatorname{Tr}(P^{-1}\hat{M}_{q})-\operatorname{Tr}(P^{-1}{M}_{q})|=|\operatorname{Tr}(P^{-1}\Delta\hat{M}_{q})|$ where $\Delta{M}_{q}=\hat{M}_{q}-M_{q}$. We first note that ${M}_{q}[i,j]={p}(j\in C_{q}(x),\tilde{y}=i)$ is not a CDF but we can define one that agrees with it for $q\in(-\infty,C)$ which is the range of interest where $C$ is a constant that bound the score function $S(x,y)$ from above. We define $\tilde{S}_{ij}(x,\tilde{y})=\begin{cases}S(x,j),&\text{if }\tilde{y}=i\\ C,&\text{if }\tilde{y}\neq i\end{cases}$ so ${M}_{q}[i,j]=p(\tilde{S}_{ij}(x,\tilde{y})\leq q)$ for $q\in(-\infty,C)$. Now from the DKW theorem, we know that if we estimate a CDF using $n$ samples then with probability at least $1-\delta$ we get a uniform bound on the error of size $\sqrt{\frac{\log(2/\delta)}{2n}}$. As we are estimating $k^{2}$ matrix elements we can use the union bound to get that with probability $1-\delta$ the $\forall i,j,q\in(-\infty,C):|\Delta{M}_{q}|\leq\sqrt{\frac{\log(2k^{2}/\delta)}{2n}}$. Now if we look at the infinity norm of $P^{-1}$, then $|(P^{-1}\Delta{M}_{q})_{i,j}|\leq||P^{-1}||_{\infty}\sqrt{\frac{\log(2k^{2}/\delta)}{2n}}$. As the trace is the sum of $k$ such matrix entries, the total bound is $\kappa_{\infty}k\sqrt{\frac{\log(2k^{2}/\delta)}{2n}}$ for $q\in(-\infty,C)$. Since we know ${F}^{c}(q)=1$ for $q\geq C$, we can set $\hat{F}^{c}(q)=1$ for $q\geq C$ and get a bound for all $q\in\mathbb{R}$. ∎
A.3 Table 1 with models trained with noisy labels
Reproducing Table 1 by replacing models publicly available with models trained with noisy labels. We combined our NACP with a noise-robust network training that estimated the noise level $\epsilon$ as part of the training [9]. As a by-product of the noisy labels training procedure, we get a noise estimation that can be used in the calibration step. Therefore the following table has three additional rows denoted by * corresponding to ACNL, NACP, and NACP (w/o $\Delta$) using $\tilde{\epsilon}$ instead of the true $\epsilon=0.2$. The estimated noise $\tilde{\epsilon}$ for each model is detailed in Table 6 and the results are shown in Table 7.
Dataset TissueMNIST OrganSMNIST OrganCMNIST OrganAMNIST PathMNIST
Estimated noise $\tilde{\epsilon}$ 0.175 0.19 0.18 0.165 0.16
Table 6: Estimated noise using for the five medical datasets, given true noise $\epsilon=0.2$.
Table 7: APS and RAPS (randomized versions) calibration results for $1\!-\!\alpha$ = 0.9 and $\epsilon=0.2$. We report the mean and the std over 1000 different splits. Bold for best result with theoretical guarantees.
Dataset CP Method rand-APS rand-RAPS
size $\downarrow$ coverage (%) size $\downarrow$ coverage (%)
TissueMNIST CP (Oracle) 3.39 $\pm$ 0.02 90.02 $\pm$ 0.21 3.39 $\pm$ 0.02 90.02 $\pm$ 0.20
Noisy-CP 5.03 $\pm$ 0.03 96.83 $\pm$ 0.11 5.04 $\pm$ 0.03 96.81 $\pm$ 0.12
NR-CP 4.07 $\pm$ 0.02 93.74 $\pm$ 0.12 4.18 $\pm$ 0.02 94.21 $\pm$ 0.12
ACNL 4.29 $\pm$ 0.04 94.63 $\pm$ 0.20 5.45 $\pm$ 0.02 97.61 $\pm$ 0.08
NACP 3.65 $\pm$ 0.03 91.67 $\pm$ 0.25 3.65 $\pm$ 0.03 91.66 $\pm$ 0.25
NACP (w/o $\Delta$) 3.39 $\pm$ 0.03 90.03 $\pm$ 0.25 3.39 $\pm$ 0.03 90.03 $\pm$ 0.25
estimated $\epsilon$ ACNL* 4.15 $\pm$ 0.04 94.08 $\pm$ 0.21 4.18 $\pm$ 0.02 94.68 $\pm$ 0.18
NACP* 3.64 $\pm$ 0.03 91.60 $\pm$ 0.24 3.64 $\pm$ 0.03 91.61 $\pm$ 0.24
NACP* (w/o $\Delta$) 3.40 $\pm$ 0.02 90.06 $\pm$ 0.23 3.40 $\pm$ 0.03 90.05 $\pm$ 0.25
OrganSMNIST CP (Oracle) 1.64 $\pm$ 0.02 90.04 $\pm$ 0.58 1.63 $\pm$ 0.02 90.02 $\pm$ 0.60
Noisy-CP 5.59 $\pm$ 0.18 99.73 $\pm$ 0.06 5.58 $\pm$ 0.19 99.74 $\pm$ 0.05
NR-CP 3.05 $\pm$ 0.06 98.23 $\pm$ 0.12 3.79 $\pm$ 0.12 99.21 $\pm$ 0.12
ACNL 2.83 $\pm$ 0.23 97.77 $\pm$ 0.48 3.20 $\pm$ 0.03 97.90 $\pm$ 0.03
NACP 1.95 $\pm$ 0.06 93.93 $\pm$ 0.64 1.94 $\pm$ 0.06 93.97 $\pm$ 0.60
NACP (w/o $\Delta$) 1.63 $\pm$ 0.03 89.94 $\pm$ 0.67 1.62 $\pm$ 0.03 90.00 $\pm$ 0.45
estimated $\epsilon$ ACNL* 2.63 $\pm$ 0.17 97.30 $\pm$ 0.53 3.08 $\pm$ 0.04 97.62 $\pm$ 0.05
NACP* 1.94 $\pm$ 0.05 93.88 $\pm$ 0.54 1.93 $\pm$ 0.05 93.87 $\pm$ 0.58
NACP* (w/o $\Delta$) 1.63 $\pm$ 0.03 89.86 $\pm$ 0.69 1.62 $\pm$ 0.03 89.86 $\pm$ 0.59
OrganCMNIST CP (Oracle) 1.22 $\pm$ 0.02 90.08 $\pm$ 0.52 1.22 $\pm$ 0.01 90.01 $\pm$ 0.66
Noisy-CP 5.50 $\pm$ 0.20 99.85 $\pm$ 0.05 5.58 $\pm$ 0.19 99.69 $\pm$ 0.05
NR-CP 2.15 $\pm$ 0.05 98.5 $\pm$ 0.16 3.42 $\pm$ 0.31 99.57 $\pm$ 0.08
ACNL 1.84 $\pm$ 0.12 97.49 $\pm$ 0.55 1.76 $\pm$ 0.10 97.41 $\pm$ 0.50
NACP 1.40 $\pm$ 0.03 94.00 $\pm$ 0.59 1.40 $\pm$ 0.03 93.97 $\pm$ 0.63
NACP (w/o $\Delta$) 1.21 $\pm$ 0.02 90.03 $\pm$ 0.62 1.24 $\pm$ 0.02 90.11 $\pm$ 0.65
estimated $\epsilon$ ACNL* 1.66 $\pm$ 0.08 96.50 $\pm$ 0.59 1.51 $\pm$ 0.05 95.80 $\pm$ 0.51
NACP* 1.39 $\pm$ 0.03 93.70 $\pm$ 0.61 1.38 $\pm$ 0.03 93.61 $\pm$ 0.54
NACP* (w/o $\Delta$) 1.21 $\pm$ 0.02 89.71 $\pm$ 0.65 1.21 $\pm$ 0.02 89.70 $\pm$ 0.76
OrganAMNIST CP (Oracle) 1.11 $\pm$ 0.01 89.98 $\pm$ 0.38 1.11 $\pm$ 0.01 90.00 $\pm$ 0.36
Noisy-CP 5.48 $\pm$ 0.12 99.87 $\pm$ 0.03 5.52 $\pm$ 0.11 99.85 $\pm$ 0.03
NR-CP 1.94 $\pm$ 0.03 98.59 $\pm$ 0.08 3.49 $\pm$ 0.19 99.65 $\pm$ 0.06
ACNL 1.35 $\pm$ 0.03 95.54 $\pm$ 0.42 1.56 $\pm$ 0.04 96.54 $\pm$ 0.39
NACP 1.19 $\pm$ 0.01 92.53 $\pm$ 0.44 1.19 $\pm$ 0.01 92.51 $\pm$ 0.37
NACP (w/o $\Delta$) 1.11 $\pm$ 0.01 89.75 $\pm$ 0.48 1.11 $\pm$ 0.01 89.74 $\pm$ 0.43
estimated $\epsilon$ ACNL* 1.28 $\pm$ 0.02 94.39 $\pm$ 0.45 1.32 $\pm$ 0.03 94.52 $\pm$ 0.40
NACP* 1.19 $\pm$ 0.01 92.33 $\pm$ 0.35 1.18 $\pm$ 0.01 92.27 $\pm$ 0.40
NACP* (w/o $\Delta$) 1.11 $\pm$ 0.01 89.78 $\pm$ 0.39 1.11 $\pm$ 0.01 89.74 $\pm$ 0.42
PathMNIST CP (Oracle) 1.10 $\pm$ 0.02 90.03 $\pm$ 0.47 1.03 $\pm$ 0.01 90.00 $\pm$ 0.45
Noisy-CP 4.81 $\pm$ 0.15 99.98 $\pm$ 0.01 4.89 $\pm$ 0.18 99.95 $\pm$ 0.03
NR-CP 1.46 $\pm$ 0.03 98.33 $\pm$ 0.19 1.46 $\pm$ 0.05 99.20 $\pm$ 0.11
ACNL 1.25 $\pm$ 0.05 95.08 $\pm$ 0.40 1.32 $\pm$ 0.06 95.30 $\pm$ 0.51
NACP 1.14 $\pm$ 0.03 93.42 $\pm$ 0.40 1.12 $\pm$ 0.01 93.48 $\pm$ 0.50
NACP (w/o $\Delta$) 1.10 $\pm$ 0.02 90.08 $\pm$ 0.30 1.04 $\pm$ 0.02 90.14 $\pm$ 0.53
estimated $\epsilon$ ACNL* 1.19 $\pm$ 0.02 93.81 $\pm$ 0.47 1.23 $\pm$ 0.03 94.00 $\pm$ 0.31
NACP* 1.17 $\pm$ 0.02 93.08 $\pm$ 0.47 1.10 $\pm$ 0.01 93.10 $\pm$ 0.45
NACP* (w/o $\Delta$) 1.10 $\pm$ 0.02 90.09 $\pm$ 0.50 1.03 $\pm$ 0.01 90.07 $\pm$ 0.49
A.4 ImageNet with different calibration set size
In the following experiment, we test the performance of various conformal prediction methods under noisy labels as a function of the calibration set size on the ImageNet dataset. Figure 3 shows the mean size and coverage as a function of the calibration set size. In addition, the correction term $\Delta$ is depicted for ImageNet for each calibration set size. Results show that even with as little as 2500 images that correspond to 2.5 images per class the calibration results are almost on par with the oracle calibration given clean labels.
(a)                                     (b)                                     (c)
Figure 3: Noisy labels conformal prediction on ImageNet with different calibration set sizes. (a) Mean size (b) Coverage (%), and (c) Correction terms $\Delta$ as a function of calibration set size.
A.5 NACP Agnostic to Different model architectures
Conformal prediction in general and our method NACP specifically has no assumption and is agnostic to the underlying model architecture. In the following section, we verify that by experimenting with ImageNet across different model architectures. Table 8 presents the results of applying conformal prediction with and without noisy labels on ResNet18, ResNet50, DenseNet121, ViT-B16 (Vision transformer).
@@ A.5 NACP Agnostic to Different model architectures
NACP (w/o $\Delta$)   16.7 $\pm$ 0.51   90.0 $\pm$ 0.34 13.9 $\pm$ 0.47 90.0 $\pm$ 0.37 12.0 $\pm$ 0.38 90.0 $\pm$ 0.34 10.7 $\pm$ 0.55 90.0 $\pm$ 0.35
rand-RAPS CP (Oracle) 6.3 $\pm$ 0.06 90.0 $\pm$ 0.27 4.5 $\pm$ 0.05 89.9 $\pm$ 0.29 4.7 $\pm$ 0.06 90.0 $\pm$ 0.26 2.6 $\pm$ 0.04 90.0 $\pm$ 0.25
@@
NACP (w/o $\Delta$)   3.6 $\pm$ 0.14   90.0 $\pm$ 0.38 2.1 $\pm$ 0.06 90.0 $\pm$ 0.38 2.4 $\pm$ 0.07 90.0 $\pm$ 0.34 1.5 $\pm$ 0.03 90.0 $\pm$ 0.35

### v2

@@ 5 Experiments
Table 2: Finite sample correction terms $\Delta$ of NACP, ACNL (Sesia et al., 2024) and CRCP (Clarkson et al., 2024), for several datasets and two noise levels, $n$ is the size of the calibration set.
Dataset $n$ #classes NACP ACNL CRCP
$\epsilon=0.1$ $\epsilon=0.2$ $\epsilon=0.1$ $\epsilon=0.2$ $\epsilon=0.1$ $\epsilon=0.2$
CIFAR-10 5000 10 0.035 0.043 0.031 0.059 0.016 0.036
CIFAR-100 10000 100 0.025 0.030 0.077 0.163 0.039 0.088
TinyImagenet 5000 200 0.035 0.043 0.175 0.382 0.078 0.176
ImageNet 25000 1000 0.016 0.019 0.194 0.466 0.079 0.177
Conformal prediction results. Table 1 reports the noisy label calibration results across 3 different conformal prediction scores, HPS, APS, and RAPS for four standard publicly available datasets, CIFAR-10, CIFAR-100, Tiny-ImageNet, and ImageNet. In all cases, we used $1\!-\!\alpha=0.9$ and a noise level of $\epsilon=0.2$. The results indicate that in the case of a calibration set with noisy labels, the Noisy-CP threshold became larger to facilitate the uncertainty induced by the noisy labels. This yielded larger prediction sets and the coverage was higher than the target coverage which was set to $90\%$. We can see that NACP outperformed the ACNL, and CRCP methods for all datasets except for CIFAR-10 with fewer classes. Following Theorem 3.5, we expect the gain in performance when using NACP versus Noisy-CP to increase with the number of classes, indeed validated empirically in Table 1. Here for CIFAR-100, Tiny-ImageNet, and ImageNet the ACNL and CRCP methods failed due to the large number of classes and the relatively small number of samples per class. A direct comparison of the finite sample correction terms $\Delta$ obtained by NACP, ACNL and CRCP is shown in Table 2. Note that if $1-\alpha+\Delta>1$, the prediction set includes all the classes and thus it becomes useless. We can see in Table 2 that this is the case for ACNL and CRCP in datasets with a large number of classes when the noise level is sufficiently high.
Correction term analysis. Following the theoretical and empirical results, the effectiveness of our method and baselines can be fully explained by the correction terms $\Delta$ each method guarantees as practitioners require coverage guarantee and therefore will use $1-\alpha+\Delta$. Note that, as explained in Section 4, our finite sample coverage guarantee is different from the one provided by the baseline method. Figure 1 presents the correction term as a function of calibration set size and the number of classes. Note that the NACP curve remains exactly the same across the 3 plots. Our main contribution is grounded in the fact that NACP is not dependent on the number of classes $k$, clearly shown in plots as the number of classes grows.
Figure 1: Correction terms $\Delta$ of NACP, ACNL and CRCP as a function of the calibration set size $n$ given $\epsilon=0.2$. We show results for 3 numbers of classes, 10, 100 and 1000.
Different network architectures. Conformal prediction in general and our method NACP specifically has is agnostic to the underlying model architecture. We next experimentally verify it with ImageNet across different model architectures. Table 3 presents the results of applying conformal prediction on noisy labels using ResNet18, ResNet50, DenseNet121, ViT-B16 (Vision transformer). We can see that the NACP yields significantly improved results regardless of the network architecture while the coverage guarantee provided by ACNL and CPRC is useless. We can see that even without adding any correction term, we get the required coverage guarantee. This indicates that all the current correction terms are very conservative.
@@
RAPS CP (oracle) 6.3 $\pm$ 0.06 90.0 $\pm$ 0.27 4.5 $\pm$ 0.05 89.9 $\pm$ 0.29 4.7 $\pm$ 0.06 90.0 $\pm$ 0.26 2.6 $\pm$ 0.04 90.0 $\pm$ 0.25
@@
6 Conclusions
We presented a procedure that applies the Conformal Prediction algorithm on a calibration set with noisy labels. We first presented our method in the simpler case of a uniform noise model and then extended it to a general noise matrix. We showed that if the noise level is given, we can find the noise-free calibration threshold without access to clean data by using the noisy-label data. We showed that the finite sample coverage guarantee obtained by current methods is not effective in the case of classification tasks with a large number of classes. We proposed a different notion of coverage guarantee and derived a suitable finite sample coverage guarantee that remains effective in tasks with a large number of classes. We showed, however, that even without adding the finite sample term, noise-robust methods obtained the required coverage. This indicates that the current coverage guarantee methods are very conservative and there is room for future research to improve it and to find much tighter bounds. In this study, we focused on noise models that assume the noisy label and the input image are independent, given the true label. In a more general noise model, the label corruption process also depends on the input features.